# ISRE-$0$-$0$.00

### v2

@@ 2.2 AI Adoption
Our paper contributes to the emerging literature on AI adoption, summarized in Appendix Table LABEL:tab:aiadoption. Prior research has predominantly focused on the antecedents and consequences of adopting traditional, non-generative AI technologies such as automated guided vehicles, computer vision, and natural language processing (McElheran et al. 2024, Chen et al. 2024, Alekseeva et al. 2020). In exploring the antecedents, studies have largely adopted a behavioral perspective to investigate how technological, organizational, and environmental factors influence AI adoption decisions (Jo and Bang 2023, Prasad Agrawal 2023). These studies primarily employ qualitative methods, including surveys and interviews, to deliberate on users’ perceptions of AI (e.g., perceived ease of use and perceived usefulness) and how these perceptions affect adoption incentives (Gupta et al. 2024, Choudhury and Shamszare 2023). Regarding the consequences, it is well established that although AI adoption produces economic benefits, these benefits may be unevenly distributed. For instance, Agrawal et al. (2024) suggests that unequal access to and capabilities in utilizing non-generative AI technologies can exacerbate disparities among firms in terms of organizational performance and market dynamics. This issue is particularly concerning, as it may lead to a widening gap between firms able to harness the benefits of AI and those that cannot, resulting in significant disparities in innovation, productivity, and market competitiveness (Gans 2023).
Our study advances beyond this literature by developing a unified model framework that simultaneously addresses both antecedents and consequences of AI adoption. While prior research has made valuable contributions, the inherently dynamic and intangible nature of user perceptions has posed significant methodological challenges. Traditional approaches struggle to directly measure the underlying utility driving adoption decisions or capture the dynamic nature of the learning process. To address these limitations, we develop a customized Bayesian learning model that formally characterizes the dynamics of generative AI adoption. This structural approach complements existing behavioral studies by moving beyond static snapshots to capture the iterative nature of technology adoption and learning.
@@ 2.3 Learning and Decision-Making under Uncertainty
These models have been most commonly applied in the context of product attribute assessment, particularly in evaluating product quality (Erdem and Keane 1996). Consumers iteratively update their beliefs regarding product attributes by utilizing various information signals, including direct experience, advertising exposure, and online reviews (Lin et al. 2015, Erdem et al. 2008). The versatility of these models extends to more complex domains, such as pharmaceutical efficacy evaluation (Ching and Lim 2020), consumption of addictive products (Chen and Rao 2020), and financial investment decisions (Zhou 2021). Moreover, within the field of information systems, Bayesian learning models have been adapted to investigate diverse user behaviors, encompassing phenomena such as opinion polarization (Lu et al. 2022), application usage patterns (Zheng et al. 2020, Zhang et al. 2020), rating behaviors (Ho et al. 2017), ideation processes (Huang et al. 2014), and content generation dynamics (Ghose and Han 2011).
@@
Our research context—individuals learning about generative AI through repeated interactions that are subject to randomness—presents an ideal setting for applying structural learning models. The inherent variability in AI system outputs creates uncertainty that users must resolve through experience, making the Bayesian learning framework particularly appropriate. To our knowledge, this study is among the first to utilize a learning model to characterize heterogeneous AI adoption patterns and to quantify the learning process as a mechanism contributing to the AI divide. Furthermore, given the ongoing evolution of AI technologies, our proposed framework offers potential for longitudinal monitoring of dynamic adoption trends.
@@ 3 Data, Variable, and Exploratory Analysis
This section describes the data source and variables for the empirical testing. We conduct reduced-form analyses to explore heterogeneous adoption and usage patterns across diverse demographic cohorts. Subsequently, we underscore the imperative for developing a learning model to formally examine underlying mechanisms, based on the exploratory findings.
@@ 3.1 Data Source
This study uses a comprehensive dataset comprising detailed daily clickstream records from users, provided by a prominent, anonymized data collection firm. Participants install a clickstream collection application on their devices, which captures a record in the database for each URL visited, including the user ID, access timestamps, and the specific URL. Complementing these online clickstream activities, the firm collects various demographic attributes, such as gender and age, facilitating social-economic analyses. Furthermore, the URLs enable direct extraction of search keywords entered into search engines or specific websites accessed by users.
@@
To elucidate the adoption dynamics of ChatGPT, we analyze a dataset spanning from November 30, 2022 (ChatGPT’s release) to May 31, 2023. The six-month duration provides sufficient time window to observe meaningful patterns in user behavior, allowing us to track how individuals discover, experiment with, and potentially integrate ChatGPT into their routines. Our sample comprises 8,306 users who were active during the observation window and for whom full demographic profiles are available. Within this cohort, 1,102 individuals have interacted with ChatGPT at least once during the observation period.
3.2 Variable Construction
With the aim of investigating the existence of a digital divide, we seek to link individual differences in ChatGPT usage with their socio-demographic and socio-economic attributes. Specifically, we examine how factors such as educational attainment, age, race, gender, and occupational sector influence the AI adoption process—factors consistently highlighted in the literature on the digital divide. For instance, educational attainment has been identified as one of the most robust predictors of digital disparities (Hidalgo et al. 2020, Hsieh et al. 2011, 2008). Digital literacy has also been strongly associated with age, with younger cohorts typically demonstrating higher proficiency (Wang et al. 2024, Jackson et al. 2001). Numerous studies have further identified significant impacts of racial/ethnic variations (Boeing 2020, Fairlie 2004) and gender disparities (Van Dijk and Hacker 2003, Jackson et al. 2001) on the digital divide. Moreover, the occupational and industrial sectors play a crucial role in shaping digital skills and usage patterns (Hidalgo et al. 2020, Venkatesh et al. 2014). To operationalize these factors, we employ a series of variables, whose descriptions and summary statistics are provided in Table 1.
In addition to demographic attributes, we introduce two individual-level variables to characterize general adoption and usage patterns. Specifically, we define $GPTUsed$ to denote whether the individual has first-hand experience with ChatGPT during the observation window. Moreover, for individuals who used ChatGPT, we construct $GPTFrequency$ to measure the frequency of interactions with ChatGPT. Summary statistics for these two variables are presented in Table 1.
Table 1: Summary Statistics (Exploration)
Variable Description N Mean Std Max Min Median
GPTUsed A binary indicator of whether the user uses ChatGPT at least once. 8,306 0.14 0.34 1 0 0
GPTFrequency Average number of ChatGPT access per day since the first usage of ChatGPT. 8,306 0.17 1.27 63.13 0 0
HighEdu A binary indicator of whether the user receives education after high school. 8,306 0.46 0.50 1 0 0
Age The user’s age in integers. 8,306 45.94 14.71 98 18 44
Male A binary indicator of whether the user is male or not. 8,306 0.39 0.49 1 0 0
White A binary indicator of whether the user is a white person. 8,306 0.74 0.44 1 0 1
IT A binary indicator of whether the user works in IT-related industry. 8,306 0.04 0.20 1 0 0
Note: For users that do not use ChatGPT at all, $GPTFrequency$ is filled with zero for summary statistics.
3.3 Exploratory Analyses
Using the constructed variables, we provide exploratory evidence of the potential digital divide. First, the differences in demographic distributions between the overall sample and the ChatGPT user subset are illustrated in Figure 1. The figures suggest that younger, male, non-white, higher-educated individuals, as well as those employed in the IT industry, are more likely to get initial exposure to ChatGPT.
Figure 1: Demographic distributions for all users and active ChatGPT users.
Given that usage frequency is measurable only among users who have accessed ChatGPT, we apply a Heckman two-stage selection model to correct for potential selection bias. In the first stage, we use a probit model to capture the heterogeneity in users’ initial access to ChatGPT based on demographic characteristics. In the second stage, conditional on individuals having used GPT at least once during the observation window, we estimate the heterogeneity in their average daily access following initial exposure. We employ the two-step estimation procedure, with results presented in Table 2.
Overall, the findings indicate that younger, male, non-white, more educated individuals, and those employed in the IT industry are more likely to use ChatGPT and engage with it more frequently. These findings, with the exception of that related to race, are consistent with the previous literature on the digital divide, which has consistently identified that younger individuals (Wang et al. 2024, Hidalgo et al. 2020), males (Van Dijk and Hacker 2003), individuals with higher educational attainment (Hsieh et al. 2011, 2008), and those with more technology-related experience (Hargittai 2003) are less likely to experience digital divides in the adoption and use of ICT. However, our findings regarding race are counterintuitive, as prior literature suggests that white individuals are more likely to adopt and have access to ICT (Boeing 2020, Fairlie 2004). This inconsistency may be explained by the possibility that non-white users face language barriers prior to using ChatGPT, which might actually drive adoption as a means to overcome language-related challenges.
Table 2: Estimation Results (Heckman)
(1) (2)
Variables GPTUsed GPTFrequency
HighEdu 0.366∗∗∗ 5.638∗∗
(0.0376) (2.468)
Age --0.0241∗∗∗ --0.3734∗∗
(0.00139) (0.159)
Male 0.517∗∗∗ 7.664∗∗
(0.0368) (3.368)
White --0.0923∗∗ --1.687
(0.0403) (0.869)
IT 0.420∗∗∗ 6.311∗∗
(0.0750) (2.852)
Constant --0.429∗∗∗ --18.09∗∗
(0.0640) (8.552)
Observations 8,306 8,306
Note: Standard errors in parentheses; ∗∗∗ p$<$0.01, ∗∗ p$<$0.05, p$<$0.1.
@@ 3.4 Variable Construction for the Bayesian Learning Model
The above descriptive analyses, however, are insufficient to disentangle utility heterogeneities and learning ability heterogeneities as potential mechanisms underlying the observed adoption patterns. To move beyond static analyses and fully reveal the underlying dynamics, it is necessary to develop a structural model that identifies the learning process while simultaneously accounting for utility differences.
To achieve this objective, we detail the construction process of the variables employed in the Bayesian learning model. First, among all individuals, we select those who have used ChatGPT at least twice to capture the learning process (Ho et al. 2017). We aggregate the clickstream data on a daily basis, resulting in 183 potential observations per individual, each representing one day’s usage. Since we focus on the learning process of ChatGPT, it is appropriate to exclude a user’s records prior to their first exposure to ChatGPT, whether through reading related news or using it directly. Following these aforementioned pre-processing steps, we construct a dataset consisting of 111,371 observations for 1,102 users.
@@
Total Activities(Search + ChatGPT) $W_{itd}$ 111,371 18.40 43.18 2,658 0 5
ChatGPT Times $W_{its}$ 111,371 1.21 11.09 2,651 0 0
Search Times $W_{itd}-W_{its}$ 111,371 17.19 41.19 1,297 0 4
News Exposure $W_{itn}$ 111,371 0.01 0.44 130 0 0
@@
High-level Education $HighEdu_{i}$ 1,102 0.57 0.49 1 0 1
Age $Age_{i}$ 1,102 39.11 13.62 88 18 38
Male $Male_{i}$ 1,102 0.60 0.48 1 0 1
White $White_{i}$ 1,102 0.66 0.47 1 0 1
IT $IT_{i}$ 1,102 0.09 0.29 1 0 0
Each time an individual $i$ accesses online resources, they may use ChatGPT or traditional search engines [^3] Here, for traditional search engines, we focus on Google and Microsoft Bing, as they account for the majority of the search engine market share.. We define $W_{itd}$ as the total number of activities in a day for a user, and $W_{its}$ represents the daily usage of ChatGPT. The remainder, $(W_{itd}-W_{its})$, corresponds to the usage of traditional search engines.
Given that ChatGPT represents a novel technology, users may rely on their prior experiences with the tool to evaluate its effectiveness and make informed decisions regarding its future use (Ma et al. 2024, Russo 2024). Consequently, ChatGPT usage (measured by $W_{its}$) not only serves as an outcome of user choice but also becomes an information source for the learning process. To account for other online information sources that may influence individuals’ beliefs (Wang and Yang 2024), we also generate a variable $W_{itn}$, which quantifies the number of times a user clicks on ChatGPT-related news articles within a given week [^4] The news outlets used were retrieved from the following website as of April 2024: https://pressgazette.co.uk/media-audience-and-business-data/media_metrics/most-popular-websites-news-us-monthly-3/. The specific processes through which users utilize these variables to revise their beliefs are detailed in Section 4.2. Consistent with the exploratory analysis, we also include $HighEdu_{i}$, $Age_{i}$, $Male_{i}$, $White_{i}$, and $IT_{i}$ as the source of digital divide in the Bayesian learning model. Table 3 presents summary statistics for all these variables.

### v3

@@ 2.2 AI Adoption
Our paper contributes to the emerging literature on AI adoption, summarized in Appendix Table LABEL:tab:aiadoption. Prior research has focused on the antecedents and consequences of adopting traditional AI technologies (Wang et al. 2025, McElheran et al. 2024, Chen et al. 2024, Alekseeva et al. 2020). In exploring the antecedents, studies have adopted a behavioral perspective to investigate how technological, organizational, and environmental factors influence AI adoption decisions (Jo and Bang 2023, Prasad Agrawal 2023). The majority of studies employ behavioral methods, including surveys, interviews, and controlled experiments, to deliberate on users’ perceptions of AI (e.g., perceived ease of use and perceived usefulness) and how these perceptions affect adoption incentives (Gupta et al. 2024, Choudhury and Shamszare 2023). Regarding the consequences, it is well established that although AI adoption produces economic benefits, these benefits may be unevenly distributed. For instance, Agrawal et al. (2024) suggests that unequal access to and capabilities in utilizing non-generative AI technologies can exacerbate disparities among firms in terms of organizational performance and market dynamics. This issue is concerning, as it may lead to a widening gap between firms able to harness the benefits of AI and those that cannot, resulting in significant disparities in innovation, productivity, and market competitiveness (Gans 2023).
Our paper models generative AI adoption using real-world data and thus complements experimental evidence and survey studies on AI adoption in two ways. First, prior experimental studies on generative AI provide important causal evidence on short-run productivity and learning effects under controlled exposure, involving fixed tasks, mandated use, and a limited interaction time window (Humlum and Vestergaard 2024, Noy and Zhang 2023). In contrast, our study examines repeated, voluntary use of generative AI over an extended observational window, where users can freely experiment, reduce usage, or abandon generative AI altogether. It allows us to uncover self-reinforcing mechanisms such as belief traps, in which early negative or ambiguous experiences lead users to prematurely exit and thereby forgo further belief updating, dynamics that cannot be detected in short-horizon or forced-use experimental designs. Second, the real-world setting allows us to disentangle the learning and utility divides. In experimental settings, learning and utility divides are intertwined because usage intensity, learning opportunities, and outcomes are constrained by design rather than users’ choice. In contrast, in real-world settings, learning speed, per-use utility, and adoption decisions are determined by users’ own choices under uncertainty rather than by experimental protocols. By combining rich behavioral data with a structural Bayesian learning framework, we are able to reveal how inequality emerges dynamically when learning unfolds over time and adoption is voluntary.
@@ 2.3 Learning and Decision-Making under Uncertainty
These models have been most commonly applied in the context of product attribute assessment, particularly in evaluating product quality (Erdem and Keane 1996). Consumers iteratively update their beliefs regarding product attributes by utilizing various information signals, including direct experience, advertising exposure, and online reviews (Lin et al. 2015, Erdem et al. 2008). The versatility of these models extends to more complex domains, such as pharmaceutical efficacy evaluation (Ching and Lim 2020), consumption of addictive products (Chen and Rao 2020), and financial investment decisions (Zhou 2021). Moreover, Bayesian learning models have been adapted to investigate diverse user behaviors, encompassing phenomena such as opinion polarization (Lu et al. 2022), application usage patterns (Zheng et al. 2020, Zhang et al. 2020), rating behaviors (Ho et al. 2017), ideation processes (Huang et al. 2014), and content generation dynamics (Ghose and Han 2011).
@@
Our research context where individuals learning about generative AI through repeated interactions that are subject to randomness, presents an ideal setting for applying structural learning models. The inherent variability in AI system outputs creates uncertainty that users must resolve through experience, making the Bayesian learning framework particularly relevant. To our knowledge, this study is among the first to utilize a learning model to characterize heterogeneous AI adoption patterns and to quantify the learning process as a mechanism contributing to the AI divide. Furthermore, given the ongoing evolution of AI technologies, our proposed framework offers potential for longitudinal monitoring of dynamic adoption trends.
3 Theoretical Development
In this section, we develop a theoretical framework for generative AI divide and explain why the underlying mechanisms may differ from the divides in traditional ICT adoption. First, we highlight key technical features of generative AI that can generate distinctive behavioral patterns. Then, we theorize how social characteristics shape generative AI outcomes through AI literacy, giving rise to learning and utility divides at the capability and outcome layers even when access is held constant. Figure 1 illustrates how our study is situated in the literature on the digital divide and how these mechanisms unfold in the generative AI setting.
Figure 1: Conceptual Framework of Generative AI Divide
3.1 Key Technical Distinctions Between Generative AI and ICTs
Generative AI models operate on a generative and probabilistic architecture. It produces responses via probabilistic token generation: when a user queries, the system generates outputs by sampling from a learned conditional distribution given the prompt (Morris et al. 2023, Kojima et al. 2022). In contrast, conventional ICTs, such as personal computers, software applications, and search engines, are relatively more predictable. They help users locate and interact with information and services that exist within or outside the system (Dewan and Riggins 2005, Venkatesh et al. 2003). These technological properties introduce fundamental distinctions between generative AI and prior ICTs.
First, generative AI can generate responses that adhere to user instructions (prompts). Most generative AI systems undergo instruction tuning, which enables them to follow specific instructions and even examples embedded in the prompt (in-context learning) to shape their outputs accordingly (Liu et al. 2025, Zhao et al. 2024, Ouyang et al. 2022). Consequently, well-crafted and individualized prompts can return personalized and niche outputs that may be difficult to surface through conventional keyword-based search queries (Xu et al. 2023). Second, generative AI produces unstructured outputs through a natural-language interface, which complicates the assessment of output quality and system capability (Zhong et al. 2024). Many prior ICTs can be evaluated using standard metrics such as accuracy and speed and easily verified. However, the output quality of generative AI is often less transparent, and supporting evidence for responses may be absent or difficult to verify, increasing the need for careful evaluation (Ma et al. 2024, Choudhury and Shamszare 2023). Third, generative AI introduces greater uncertainty and signal noise. Because generation prioritizes statistical plausibility rather than factual fidelity, responses can be fluent yet incorrect or fabricated (“hallucinations”) (Xu et al. 2024). Consequently, user feedback is noisier than in traditional ICT systems. This heightened uncertainty alters how users evaluate and update beliefs about system capability over time, making learning more important yet fragile (Yan et al. 2024).
3.2 From Technical Distinctions to Generative AI Divide
Generative AI introduces inequality because variation in AI literacy shapes how users learn from and extract value from the system over time. In our context, AI literacy manifests as the ability to formulate prompts effectively, interpret outputs accurately, and adapt usage strategies under uncertainty. First, at the capability layer, the generative nature of generative AI produces unstructured outputs, and the quality of responses to open-ended questions cannot be evaluated by accuracy alone, giving rise to a learning divide. Users have to infer system capability through repeated interactions. Those with higher AI literacy are better able to integrate information across interactions and update their beliefs, enabling faster convergence toward accurate assessments of system capability (Ng et al. 2021). Second, at the outcome layer, the instruction-following and in-context learning capabilities of generative AI create a utility divide through prompt-dependent value creation. Because the output quality depends on how users formulate prompts and frame tasks, users with higher AI literacy can elicit more relevant and accurate responses, thereby obtaining greater utility per use (Annapureddy et al. 2025). Third, the stochastic and uncertain nature of generative AI amplifies heterogeneity at the capability and outcome layers. Early interactions generate noisy and potentially misleading signals, which increase dispersion in belief updating across users and heighten the risk of mislearning or premature abandonment among users with lower AI literacy (Yan et al. 2024). Meanwhile, uncertainty increases variance in realized per-use utility, as identical prompts may yield different outcomes across interactions. These dynamics generate self-reinforcing belief traps that operate within the learning and utility divides.
In our context, AI literacy is a generative AI specific extension of digital literacy that is socially produced (Celik 2023). Consistent with research on the digital divide, AI literacy is shaped by socio-demographic characteristics (e.g., age, gender, and race) and socio-economic attributes (e.g., education, English literacy, and occupation) (Wang et al. 2024, Celik 2023, Robinson and others 2020, Scheerder et al. 2017, Dewan and Riggins 2005, Van Dijk and Hacker 2003). Variation in these social characteristics might translate into systematic differences in users’ ability to interact with, interpret, and evaluate generative AI systems, which in turn generate learning and utility divides (Vassilakopoulou and Hustad 2023, Lythreatis et al. 2022). Specifically, users who are better educated, younger, male, white, more proficient in English, and employed in occupations with greater exposure to large language models are more likely to possess higher AI literacy (Bassignana et al. 2025, McElheran et al. 2024, Fairlie 2004, Jackson et al. 2001), enabling faster learning about system capabilities and higher per-use utility from interactions with generative AI (Annapureddy et al. 2025, Celik 2023). Appendix K elaborates more on the rationales.
@@ 4 Data, Variable, and Exploratory Analysis
To empirically test our research questions, we use a real-world dataset from a clickstream tracking company, which offers several advantages (Chiou and E. Tucker 2022, Chesnes et al. 2017, Zheng et al. 2012). First, it records all URLs visited by each participant over an extended period, enabling us to observe the temporal dynamics of generative AI adoption and use. Second, in addition to clickstream records, it includes measures of participants’ social characteristics, allowing us to link these characteristics to usage patterns and empirically examine the AI divide. Third, relative to laboratory settings, this real-world data measurement enables us to reveal how inequality emerges dynamically in voluntary adoption when learning unfolds over time and uncover self-reinforcing mechanisms such as belief traps.
In this section, we describe the data source and collection procedures. We then assess the representativeness of the sample. Finally, we elaborate on the variable construction and the rationale underlying our measures to support the structural estimation.
@@ 4.1 Data Source
The data tracking company (thereafter the company) advertises for and recruits participants to contribute data in exchange for compensation. Those participants are referred to as “panelists.” To enroll, panelists complete a registration form providing social characteristics information, such as gender and age. The company verifies this information prior to approval. Once approved, panelists install a tracking application on their primary devices. The application then continuously records their clickstream activity. The company conducts ongoing targeted outreach and recruitment across demographic groups to ensure that the panel remains broadly representative of the U.S. population.
@@
To study the dynamics of ChatGPT adoption, we obtain a dataset covering November 30, 2022 (the initial public release of ChatGPT) through May 31, 2023. This six-month window allows us to observe how individuals discover, experiment with, and potentially incorporate ChatGPT into their routines. Our sample includes 11,752 users who were active for at least 60 days during the observation period and have their demographic profiles verified. Among them, 1,386 users had used ChatGPT at least once. Within this cohort, 1,344 users interacted with ChatGPT at least twice, providing the repeated observations necessary to identify choice dynamics and learning (Ho et al. 2017). In addition, we obtain the November data for the full sample to support the construction of pre-treatment behavior-based variables, as described in the next subsection.
4.2 Variable Construction and Exploration
To examine the potential presence of digital divide in ChatGPT adoption, we link individual differences in ChatGPT usage to their social characteristics. Based on our theoretical development (Section 3), we examine how education, age, gender, race, occupation, and English literacy shape the AI adoption process dynamically. Education has been identified as one of the most robust predictors of digital disparities (Hidalgo et al. 2020, Hsieh et al. 2011, 2008). In addition, digital literacy has been shown to be strongly associated with age, with younger cohorts demonstrating greater proficiency about technologies (Wang et al. 2024, Jackson et al. 2001). Prior studies further document significant roles of gender disparities (Van Dijk and Hacker 2003, Jackson et al. 2001) and racial differences (Boeing 2020, Fairlie 2004) in the digital divide. Occupation plays a crucial role in shaping digital skills and usage patterns (Hidalgo et al. 2020, Venkatesh et al. 2014), and, more relevant to generative AI adoption, the exposure to LLM-relevant tasks can substantially influence individuals’ adoption behavior (Eloundou et al. 2024, Liu et al. 2023). Lastly, because LLM performance varies across languages and often favors English, we also assess the impact of users’ English literacy (Martínez-Murillo et al. 2025, Qin et al. 2025).
We then construct a set of variables to operationalize these factors. Education is captured by $College$, a binary variable equal to 1 if a user has some college education and 0 otherwise. Age is grouped into three categories (low, middle, and high) based on the sample tertile cutoffs (35 and 47 years), with the middle and high categories represented by the dummy variables $AgeMid$ and $AgeHigh$, respectively. Gender is measured by $Male$, a binary variable equal to 1 for male users and 0 otherwise. Racial background is represented by $White$, a dummy variable indicating whether the user identifies as white. To measure occupational exposure to LLMs (LLMOE), we construct the $LLMOE$ score from Eloundou et al. (2024): we match each user’s occupation to the occupations reported in Eloundou et al. (2024) via semantic similarity and then assign the $LLMOE$ value of the most similar occupation. Finally, $English$ proxies for English literacy and is measured as the share of English language queries among all search queries made prior to ChatGPT’s launch. The summary statistics are reported in Appendix L. To assess the representativeness of our dataset, we compare the distributions of social characteristics in our sample with those in the U.S. population. Appendix N presents the comparison. Overall, there is no significant difference across social characteristics of the two distributions, suggesting that our sample is reasonably representative. As reported in Appendix O, we also test the pairwise correlation of each pair of social characteristic variables, and all correlations are smaller than 0.2, which indicates that multicollinearity is unlikely to be a concern in our setting.
We next present exploratory evidence of potential divides in generative AI adoption and usage. First, Appendix M compares the demographic distributions of the full sample with those of ChatGPT users. The comparisons suggest that adoption of ChatGPT is more likely among younger individuals, men, non-white individuals, those with some college education, those with lower English literacy, and those with higher LLMOE. Second, our reduced form analyses reinforce these patterns. Specifically, within the repeat ChatGPT user subset, we regress the average daily ChatGPT usage frequency, $GPTFrequency$, on the same set of social characteristics. As Appendix P reports, the groups highlighted in the graphical evidence (younger, male, with some college education, with lower English literacy, and with higher LLMOE) also use ChatGPT more frequently. Taken together, these results suggest the presence of a generative AI divide, which further underscores the importance of our research questions.
@@ 4.3 Variable Construction for the Bayesian Learning Model
To further disentangle utility heterogeneities and learning ability heterogeneities as underlying mechanisms of the adoption patterns, we develop a structural model that identifies the dynamic learning process while accounting for utility differences.
Modeling generative AI adoption requires an outside option against which to benchmark users’ decisions. We use search engine activity as this option, which is directly observed in our dataset. Because ChatGPT is trained on a vast corpus to answer a wide range of questions (Morris et al. 2023), it can serve as an alternative tool for accessing information. Our choice is also backed up by a growing literature that compares user search behavior in ChatGPT versus traditional search engines across different scenarios (Sun et al. 2024, Karunaratne and Adesina 2023, Xu et al. 2023). Leveraging users’ time-varying usage of ChatGPT relative to search engines, we can trace perception dynamics and disentangle the underlying mechanisms.
We construct the variables used in the Bayesian learning model in several steps. From the full sample, we retain 1,344 users who used ChatGPT at least twice, and the repeated observations allow us to observe belief updating (Ho et al. 2017). We then aggregate each user’s clickstream data at the daily level, yielding up to at most 183 observations per user, with each observation corresponding to a single day. Because our focus is on the dynamic learning process following the initial exposure to ChatGPT, we exclude records prior to a user’s first exposure to ChatGPT, whether that exposure occurs through consuming related news or through direct use. After preprocessing, the final dataset contains 134,022 observations from 1,344 users. Table 1 reports summary statistics for all variables.
@@
Total Activities(Search + ChatGPT) $W_{itd}$ 134,022 16.49 40.89 2,658 0 3
ChatGPT Times $W_{its}$ 134,022 1.18 10.49 2,651 0 0
Search Times $W_{itd}-W_{its}$ 134,022 15.31 30.98 1,297 0 2
News Exposure $W_{itn}$ 134,022 0.01 0.40 130 0 0
@@
College Education $College_{i}$ 1,344 0.56 0.50 1 0 1
Age - High Group(¿47) $AgeHigh_{i}$ 1,344 0.23 0.42 1 0 0
Age - Middle Group(35 ~ 47) $AgeMid_{i}$ 1,344 0.29 0.45 1 0 0
Male $Male_{i}$ 1,344 0.60 0.49 1 0 1
White $White_{i}$ 1,344 0.65 0.48 1 0 1
LLM Occupational Exposure $LLMOE_{i}$ 1,344 0.12 0.10 0.79 0 0.11
English Literacy $English_{i}$ 1,344 0.79 0.15 1 0 0.83
Each time a user $i$ accesses online information resources, they may use either ChatGPT or traditional search engines [^3] For traditional search engines, we focus on Google and Microsoft Bing, which together account for the majority of U.S. search engine market share.. Let $W_{itd}$ denote the total number of information-seeking activities for user $i$ on day $t$, and let $W_{its}$ represent the number of ChatGPT activities on that day [^4] One ChatGPT activity is defined as a user visiting the ChatGPT website (i.e., opening a ChatGPT page). We treat an activity as ending when the user leaves the ChatGPT page to visit other pages, or when the user opens a new ChatGPT chat that changes the URL.. The remaining activities, $W_{itd}-W_{its}$, correspond to usage of traditional search engines. Moreover, to account for other online information sources that may influence beliefs (Wang and Yang 2024), we construct $W_{itn}$ to capture the number of times user $i$ clicks on ChatGPT-related news articles on day tt [^5] The set of news outlets was compiled from the following website as of April 2024: https://pressgazette.co.uk/media-audience-and-business-data/media_metrics/most-popular-websites-news-us-monthly-3/.. Finally, consistent with our exploratory analysis, we include $College_{i}$, $AgeHigh_{i}$, $AgeMid_{i}$, $Male_{i}$, $White_{i}$, $LLMOE_{i}$, and $English_{i}$ as sources of heterogeneity underlying the generative AI divide in the Bayesian learning model.